OpenStudy (anonymous):

what does -1/2(tan(x) limit -pi/3to pi/3 equal to

OpenStudy (anonymous):

i am having trouble understanding your statement. Can you state the problem as a sentence?

OpenStudy (anonymous):

the problem is 1/2 (tan x) the limit is -pi/3 to pi/3 (you have to plug these values into x) i didn't get the right answer, i want to c what the problem is

OpenStudy (anonymous):

is this calculus?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\lim_{x\rightarrow -\frac{\pi}{3}}\frac{1}{2}\tan x\]

OpenStudy (anonymous):

is that the question?

OpenStudy (anonymous):

no i found the integral of the problem which came out to be (1/2tanx) the limit is -pi/3 to pi/3. could u plug it in and tell me what u get

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

you should state your question more clearly

OpenStudy (anonymous):

\[\frac{\sqrt{3}}{2}\]

OpenStudy (anonymous):

how did u get a positive number

OpenStudy (anonymous):

was the definite integral \[\int^{\frac{\pi}{3}}_{-\frac{\pi}{3}}\frac{1}{2} \sec^2 x dx\]

OpenStudy (anonymous):

at any rate \[\frac{1}{2}\left\{\tan(\frac{\pi}{3})-\tan(-\frac{\pi}{3})\right\}\]

OpenStudy (anonymous):

\[=\frac{1}{2}\left\{\frac{\sqrt{3}}{2}-(-\frac{\sqrt{3}}{2})\right\}\]

OpenStudy (anonymous):

\[=\frac{1}{2}\frac{\sqrt{3}}{4}\]

OpenStudy (anonymous):

sorry that was wrong

OpenStudy (anonymous):

latex error

OpenStudy (anonymous):

\[\frac{1}{2} 2\sqrt{3}\]

OpenStudy (anonymous):

did it again

OpenStudy (anonymous):

so -pi/3= sqrt3/2

OpenStudy (anonymous):

no

OpenStudy (anonymous):

i mean tan of that angle

OpenStudy (anonymous):

but \[\tan(-\frac{\pi}{3})=-\frac{\sqrt{3}}{2}\]

OpenStudy (anonymous):

got ya, but no

OpenStudy (anonymous):

man

OpenStudy (anonymous):

tangent is negative in the fourth quadrant

OpenStudy (anonymous):

look at the tangent graph or the unit circle to comfirm

OpenStudy (anonymous):

k ty